MOBY-L264 experiment with MOBYrefresh - BS01cfg007 MOBY262 postcals

 

Problems seen (and solved) and good things:
  1. Date/Time not set on BS - Art gave code to fix this
  2. Camera hanges when changing shutters - This was a problem from MOBY262, still need to fix this and check that it is not a problem for BS04 on MOBY264

Log Sheets

Photos

All the raw data files and their images , Table of KEYWORDS variable
Page Number
Link
Description
Date
1.01
Day 01 - 29 Sep Raw Data , Table
DAY01 - Mike testing aquisition via linux on BS - we thick this is a test file from Art an no Mike's data (not processed)
03 Oct 2017
1.02
Day 02 - 30 Sep Raw Data , Table
DAY02 - Retrying data collection and some Es track 1 data
03 Oct 2017
1.03
Day 03 - 2 Oct Raw Data , Table
DAY03 - LED Flashlight test scans on LuTop / Trk#9
03 Oct 2017
1.04
Day 04 - 3 Oct Raw Data , Table
DAY04 - LED Flashlight test scans on LuTop
05 Oct 2017
1.05
Day 05 - 5 Oct Raw Data , Table
DAY05 - Wavecal on LuBot
05 Oct 2017
1.06
Day 06 - 6 Oct Raw Data , Table
DAY06 - Es Post-cals using F600(C3)
06 Oct 2017
1.07
Day 07 - 10 Oct Raw Data , Table
DAY07 - Track 13 int time/exposure time calibrations
10 Oct 2017
1.08
Day 08 - 13 Oct Raw Data , Table
DAY08 - Track 1 lamp and solar wavecal
13 Oct 2017
1.09
Day 09 - 19 Oct Raw Data , Table
DAY09 - Track 9, 11 and 13 Lus post-deployment cal
13 Oct 2017
 
Day02-04 - Looking at track defintions for these days
2.01 findtracks Find tracks from day 2-4 05 Oct 2017
Day05 - LuBot wave cal with Ar, HgA, Ne, Kr, Xe lamps
6.01 Rough wavecal - T13 Rough wavecal using 5 lamps (Ar, HgA, Ne, Kr, Xe lamps) 05 Oct 2017
Day06 - Es Post-cals using F600(C3)
7.01 Es rsp Es postcal 06 Oct 2017
Day07 - Track 13 int time/exposure time calibrations
7.01 int time cal My intial look at Mike int cal data 10 Oct 2017
Day08 - Track 1 lamp and solar wavecal
8.01 Rough wavecal - T1 Rough wavecal using 5 lamps (Ar, HgA, Ne, Kr, Xe lamps) 13 Oct 2017
8.02 Fraun wavecal - T1 Rough wavecal using solar data and fraunhofer lines 13 Oct 2017
8.03 comp wavecals Comparing the two wavecals 23 Mar 2018
8.04 Es UV bump Looking at the Es UV bump 18 Oct 2017
Day09 - Track 9, 11 and 13 Lu system resp postcal
9.01 Lu rsp Lu rsps 23 Oct 2017
       

 

pre-experiment EMAIL from Mike

Hi Art,

I got the first .fits file from the BS Thru night, but I had some problems with running the camera TEC @ -40degC  - there was a constant beep tone coming from inside the BS house after the temp got down ~-40, and then the cooler stopped (pwr sup dropped from ~5 to ~2Adc).
But TEC was OK on Thur @ -20degC.  I was only pointing a fan at the BS cooling fins Thur nite. Mark said the max temp differential for the cooler is 20degC, and I needed to run an AC on the BS. Fri night I pointed an AC at the cooling fins and the temp was able to stabilize at -40degC.

The one signal .fits I collected Thur looked like a dark scan, and the "READTIME = 2e-7" - I think READTIME == Horizontal shift, which I thought I set to 3MHz, 1/3e6 = 3.3333e-7, so I think the HShift was really 5MHz, 1/5e6 = 2e7...

Fri night I seem to have gotten the HShift set OK, since "READTIME = 3.3333e-7".

Both nights the .fits 'DATE' = 2011-01-01", but the time stamp in "DATE" string was changing, ex. '2011-01-01T01:13:46',  '2011-01-01T01:31:46', '2011-01-01T01:52:28'. I was going to look into how to read-&-set date/time on the Falcon today.

Friday night, I got reasonable sig & drk ADU levels @ 10 & 30 sec exposure times, from the Es collector via an incandescent flashlight source.

But, whenever I tried to open or close the external shutter, the camera TEC stopped. I.E. the pwr sup dropped to ~2Adc.  When I checked the "f. Cooling Status" it said "Cooler is ON" but the returned "Temperature" kept increasing to ambient, and I had to Stop & Set & Start Cooling to get cooling on again.  I'm going to ax Mark today to check if I have the pwr sup set correctly - it never did go into over-voltage shutdown...

When I tried to collect a background image via "d. Set Shutter for Dark Image", then "a. Start Acquisition", there was no response, then after another "a" I got an OK "Mean = ...; Min = ..., Max = ..." response, but no .bmp nor .fits files were created, then I think the program hung and I had to ^C out & restart ./cooling. I will try another bac scan today and look more closely at this.

I did save PuTTY logs, but the window running ./cooling is kinda messy!
I was thinking two improvements might be
1.) maybe only display the full "Menu" at startup ?
2.) maybe print a time-stamp with the response to commands ?

I am going to writeup 2x log sheets for Thru + Fri & send logs + data to Stephanie for webpage display.

I think my to-do list for today is:
1) Look into how to read-&-set date/time on the Falcon (ex. is DATE = 2011-01-01?).
2) Check if time stamp .fits "DATE" string matches the Falcon time.
3) Check if I have the pwr sup set correctly - for shutter & cooler operation together.
4) Try to collect another bac scan.

Cheers, MF


DAY 1/2 EMAIL from Mike

Hi Steph,
There are some new day01 & day02 BS01 data for you and the MOBY-refresh web pages at
/ftp1/Mike/L264/bscal/ = s,d 20170929,30 001,2.fits

with log sheets at
/ftp1/Mike/L264/bscal/doc

Art's software collects a .fits file and a .bmp file.  I doubt I'll be saving the .bmp files,
unless you think they are useful.  The .bmp files are at
/ftp1/Mike/L264/bscal/bmp

I think the Day01 data - s20170929001.fits & .bmp is bogus.  I now believe it was a
test file that Art collected earlier in the day when he updated Linux software on the BS,
not the test file that I thought I was collecting later that night...

I don't think there is much to be done with the day02 data, except maybe verify the
.fits file settings in your "table" page, and maybe look at the Es track edges from
the s20170930001 & 002.fits files (?).

I think maybe I'll try to put a flashlight on the 3x Lu collectors tonight to verify
I can open the proper (external) shutters and get a signal OK - maybe those will
also be useful to check track edges also ?

Oh, that's one thing I forgot to put on the log sheets, the BS01 Es is at track #1!
Cheers, MF

DAY 1/2 reply EMAIL from Art

Hi Mike

The BSG sets its time from the modem (or the internet). If no modem connected then it will be unable to get the time automatically and will use the default startup time (jan 1, 2011, 0Z). You can set the date manually with a command like this:

sudo date -u MMDDhhmmCCYY.ss

where MM= month DD = day hh = hour etc. You can type “man date” to get a summary of the command You need the sudo to elevate your user permission to be able to do it. It should ask you for the password, it’s the same as the login password.

I have found that if there are any hiccups at all (e.g. something like the power supply issue you mentioned) you must at a minimum quit and restart the cooling program to allow it to reinitialize the camera. Possibly might even require power cycling, which means rebooting the whole thing. Maybe that is what happened when you got the strange READTIME.

I remember now: the external shutter does cause a problem with this BSG at least (maybe all of them?) You need to quit the program and reinitialize whenever changing the external shutter. We ran into this issue during the last deployment and we stopped doing the darks (external shutter closed and internal shutter open) and only did the backgrounds (external shutter open and internal shutter closed). I’m sure that’s the problem you’re having. We should discuss with Mark how to fix this (and test if it is a problem on the currently deployed BSG…)

The cooling menu is messy but not to fear the new acquisition program will be much better .. For the time being, have you tried stretching the putty window (you can make it taller) At least enough to see the whole menu at at time. That helps somewhat.

Yes the image16 files are ones I made while testing (with “A”). The image (no “16”) files are ones you made (with “a’). It’s OK to remove all of the image* files to start with a blank slate.

Art Gleason

DAY 1/2 reply EMAIL from Mike

Hi Art,
Rodger the date setting.  I tried to figure out the cryptic Linux help
last night but was unsuccessful!  I did not try "sudo", nor the "-u" switch.

Rodger the possible "cooling" prog restart after changing shutters.
Last nite it seemed to work OK though - no cooler dropouts.  I did make one
cfg change to the pwr sup, but I think it only controlled the constant-current mode,
which was not in use...

One thing I forgot to do last night was re-try a background image.
On Thursday night the one bac I tried first made no .fits file,
then froze the cooling prog.  I'll try bac files tonight.
MF

DAY 3 EMAIL from Mike

Hi STeph,

There are day03 .fits at /ftp1/Mike/L264/bs1cal/day03 with (all) logsheets renamed at /ftp1/Mike/L264/bs1cal/doc, plus day03 PuTTY logs also at /ftp1/Mike/L264/bs1cal/doc.

Day03 scans were all intended to be via LuTop / Trk#9 via LED FlashLight source, but sig #1 & #2 had the FLite on the (wrong) MOS LuTop collector = bogus scans, and sig #3:5 had the FLite on the BS LuTop collector.

It looks like all day03 scans had Horiz Shift Speed = 50kHz ( i.e. 'READTIME' [2e-005] in the fitsinfo() )
but I thought I set Horiz Shift Speed = 3MHz ( i.e. READTIME should have been 1/3e6 == 3.3333e-7  ) ...which I'll be looking at again tonight...

MF

DAY 4 EMAIL from Mike

Hi Steph,
There are a few .fits files for you from day04 at /ftp1/Mike/L264/bs1cal/day04, plus 3x log sheets & 1x new *putty.log at /ftp1/Mike/L264/bs1cal/doc.

Things went much smoother tonight.  I was able to set the Linux system clock, so the Date/Time in the .fits files are OK.  I looked at the log from yesterday and found that I had set the Horiz Shift Speed (aka READTIME) incorrectly last night (option 3 selected 0.05MHz, while setting 1 = 3MHz), so I corrected
that tonight.  The camera "background" scan worked just fein.  And we got good LED flashlight scans from 3x Lu collectors OK - the 1st signal file = s201701004001.fits was a bogus LuMid scan because I did not open the external shutter properly.  And the new power supply setup maybe cured the problem of the CCD cooler shutting down when shutter(s) were open/closed.

We might have an OK set now of Es & Lu lite scans to compare track position versus the last in-water scans ?

It might be getting near time to take some real post-cal data. I'm thinking to start with wave cal lamp scans tomorrow...
Cheers, MF


DAY 5 EMAIL from Mike

Hi Steph,

I got some wave lamp data tonight from the BS01.
Just 1lt + 1dk scans per lamp, at 5x lamps = Ar, HgA, Ne, Kr, Xe.
The Ne saturated at 0.05 sec exposure time (couldn't really go any faster than that)...
10x new *20171005*.fits files are at /ftp1/Mike/L264/bs1cal/day05/, and
1x log sheet scan + 2x 20171005*_putty.log files are at /ftp1/Mike/L264/bs1cal/doc/.
Talk to you at 10AM tomorrow! 

MF

DAY 6 EMAIL from Mike

HI SF,
Day6 BS cal data at /ftp1/Mike/L264/bs1cal/day06
are the Post-MOBY262 BS01cfg007 Es Post-cals =
2xDk,5xLt,2xDk scans,
which used the same F600(C3) cal lamP as at Pre-cal.

At /ftp1/Mike/L264/bs1cal/doc
are 1x log sheet + 2x *Putty.log files from day6,
and a writeuP of BS acquisition notes culled from Art emails =
00_BS_acq_notes.txt

There will be no new BS data collected tonight.  The OL425 may have
comPleted its' cal run at FASCAL and might be back in HI next week,
so the BS Lu Post-cals can hoPefully use that...
Aloha, MF

DAY 7 EMAIL from Mike

Hi Steph,
I got us some more BS01 data tonight!
Day07 would be the post-MOBY262 BS01 exposure time cal.

There are 27x pairs of sig & dark .fits file/scans at /ftp1/Mike/L264/bs1cal/day07,
and there is 1x new log sheet at /ftp1/Mike/L264/bs1cal/doc
(i.e. moby262poscal_BS1_log_07=day07.jpg)
plus, as an extra bonus, I re-uploaded the day06 logsheet
(i.e. moby262poscal_BS1_log_06=day06.jpg)
after fixing it so that it says "day06" at the top.

Tomorrow, I was going to try running thru my version of processing these
day07 data, and see what I get compared to the pre-cal on 12-Feb-2017,
and see what I get compared vs your processing.
Thanks eh, MF

DAY 8 EMAIL from Mike

Hi STeph,
Your NAS has for you some more data from the BS01 today =
Es scans of sunlight & Es scans of 5x Oriel wave lamps,
plus 2x log sheets  &  2x *putty.log's  &  5x photos of the Solar setup.

To separate the Sun scans from the lamp scans, I named the
Sun = s,d201710130xx.fits, and I named the lamps = s,d201710131xx.fits,
i.e. sequence numbers 0xx & 1xx.
Let me know if this does or does not work for you.
MF

DAY 9 EMAIL from Mike

Hi Steph,
I am FTPing to "our" NAS the day09 data for post-MOBY262 BS01 Lu cals
@ /ftp1/Mike/L264/
The LuBot scans are *1xx.fits, LuMid are *2xx.fits, LuTop are *3xx.fits.
The scanning sequence was 2xDk, 5xLt, 2xDk scans for each sensor,
so there are dark scans = d*01:04.fits, and lite = s*01:05.fits.
There are also 1x log sheet scan and 2x *_putty.log files.

These Lu cals were via the new Lamp#10 in the OL425,
which had an average photodiode monitor reading of 23.12 fL.
The PDmon readings during BS01 scans (see logsheet) were between 23.12 & 23.35 fL.

Here is the file for the OL425(L10) begin-of-life radiance cal
= NIST FASCAL + uncertainty + my BBfits:  <attached: OL425L10.txt>
MF